Figure 1: P_{n}-phase Event 054: Scaled data, y_{k}, k = 1, ..., q, where q = 5 and n = 220.

Figure 2: P_{n}-phase Event 054: Posterior estimates of the pulse sequence, â_{j}, j = 1, ..., m, where m = 40.

Figure 3: P_{n}-phase Event 054: Posterior estimates of the underlying signal-path effects, ŝ_{k}, on each channel.

Figure 4: P_{n}-phase Event 054: Solid: Scaled data, y_{k}. Dots: Posterior estimates, ŷ_{k}, shifted one point to the left.

Figure 5: P_{n}-phase Event 054: Last 5,000 iterations in the Markov chains produced for the peaks in â_{j}, j = a, ..., m.

Figure 6: P_{n}-phase Event 054: Last 5,000 iterations in the Markov chains produced for the peaks in â_{j}, j = a, ..., m.

Figure 7: P_{n}-phase Event 054: Posteriors of a_{j}, j = 1, ..., m.

Figure 8: P_{n}-phase Event 054: Markov chains for ϕ.

Figure 9: P_{n}-phase Event 054: Markov chains for η and τ.

Figure 10: P_{n}-phase Event 054: Posteriors of ϕ_{1}, ϕ_{2}, ϕ_{3}, η, and τ.

Figure 11: P_{g}-phase Event 054: Scaled data, y_{k}(t), k = 1, ..., q, t = 1, ..., n, where q = 5 and n = 220.

Figure 12: P_{g}-phase Event 054: Posterior estimates of the pulse sequence, â_{j}, j = 1, ..., m, where m = 40.

Figure 13: S_{n}-phase Event 054: Scaled data, y_{k}(t), k = 1, ..., q, t = 1, ..., n, where q = 5 and n = 256.

Figure 14: S_{n}-phase Event 054: Posterior estimates of the pulse sequence, â_{j}, j = 1, ..., m, where m = 40.

Figure 15: L_{g}-phase Event 054: Scaled data, y_{k}(t), k = 1, ..., q, t = 1, ..., n, where q = 5 and n = 512.

Figure 16: L_{g}-phase Event 054: Posterior estimates of the pulse sequence, â_{j}, j = 1, ..., m, where m = 50.

Figure 17: Event 054: Posterior estimates of the pulse sequence from each phase.